OpenStudy (anonymous):

Slove for x cos(34degrees)=50/x

OpenStudy (dumbcow):

multiply both sides by x x^2 cos(34) = 50 x^2 = 50/cos(34) x = +-sqrt (50/cos(34))

OpenStudy (anonymous):

kinda confused..

OpenStudy (dumbcow):

which step

OpenStudy (dumbcow):

oh wait i think i got the problem wrong sorry cos(34) = 50/x right?

OpenStudy (dumbcow):

x = 50/cos(34)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay so i did that on my calculator and I got 60.31089743. is that right?

OpenStudy (dumbcow):

correct